OpenStudy (anonymous):

-15=5r+12-2r+6 ( Show Work)

Directrix (directrix):

-15=5r+12-2r+6 -15 = 5r -2r + 12 + 6 ... so

Directrix (directrix):

-15 = 3r + 18 3r + 18 = -15 so ..

Directrix (directrix):

3r = -15 - 18 3r = -33 (1/3) 3r = (1/3) (-33) r = - 11 (check for errors, LovePiko) :)

OpenStudy (anonymous):

All Correct because I Got The Same Thing! :D

Directrix (directrix):

Congrats and thanks for the medal. :)